During opening, we again saw the secondary E axis Galil fail to stop on a full step, despite further work on the Galil by Jon B. and Wendell this afternoon, after which Steph and Russell exercised the secondary without incident. The first time we homed the axis, tcc_prim reported a 148 microstep error. We immediately saw the same problem as soon as we re-enabled the collimate process again. This time, homing showed a 1 microstep error. We did not see the problem again during the night. In pursuit of the source of the difference between the two pointing models taken this run, Robert and Steph verified that the same fiducial tables were used, and Craig and Russell eliminated axisDtime errors and problems with the earth orientation data block as possible causes. The mystery continues. We opened at 20:30 under partly cloudy skies. After fixing the problem with the secondary, we attempted spectroscopy, starting with the pointing model of last night. Plate 627, cartridge 1. No stars in guide fibers after initial gotoField. The first fk5InFiber produced bizarre behavior. The field was at (RA, Dec) = (313, -6). The track to the fk5 aimed at (248, 47)!!! This occured between getclock times 997763607, and 8997763798. Another fk5InFiber sent the telescope to a star a few degrees from the field. However, we could not find a star within a 30 arcsec radius of a spiralPattern search. We tried reverting to the previous pointing model without success. We decided to try a different cartridge. Plate 636, cartridge 3. With the _latest_ (20-point) pointing model, we found an fk5 star at (deltaRA, deltaDec) = (12, 2) from the aimed position, and quickly got on the field. Cart 1, no star with either model. gettime from 997763607, 997763798 Cart 3. field at ra=313, dec=-6. fk5infiber went to 260 +47? second fk5infiber went to nearby star. Offset ra+12, dec+2 to center star, then we got guide stars. Spectro 1 fill faults at 22:19 did not correspond to actual warming. Closed by high humdity at 00:05 MDT. i+trim2 reported -6.68V for CCD61. Watched seemed to report this about midnight (while we were closing), but a plot showed this had been the value since 10am this morning.